Find the centres of the two spheres which touch the plane 4x+3y=47 at the point (8,5,4) and the sphere x^2+y^2+z^2=1


1
Expert's answer
2021-03-04T15:23:06-0500

Solution.

Find the equation of normal to the plane:


"\frac{x-8}{4}=\frac{y-5}{3}=\frac{z-4}{0}=k,"

where "k" is some number.

Then

"C(4k+8;3k+5;4)-"

center of required spheres.

Radius of required spheres:


"R=\sqrt{(4k+8-8)^2+(3k+5-5)^2+(4-4)^2}=\newline
=\sqrt{25k^2}=5k."

Sphere "x^2+y^2+z^2=1" has center "(0;0;0)" and radius "1."

We will have equation:

"(1+5k)^2=(5k)^2+105,
\newline \text{ where } 105"is distance between (0;0;0) and (8;5;4).

From here

"1+10k+25k^2=25k^2+105,
\newline
k=10.4"

So, "(49.6;36.2;4)" and "(-33.6;-26.2;4)" are the centres of spheres.

Answer. "(-33.6;-26.2;4), (49.6;36.2;4)."